OpenStudy (anonymous):

A total of 636 tickets were sold for the school play. They were either adult tickets or student tickets. There were 64 fewer student tickets sold than adult tickets. How many adult tickets were sold?

OpenStudy (lucaz):

S=students, A=adults 64 fewer student tickets -> A-64 636 = A+S 636 = A+(A-64)

OpenStudy (lucaz):

solving for A you get 350 adult tickets, and S=A-64=350-64=286 student tickets

OpenStudy (anonymous):

your awesome thank you